OpenStudy (anonymous):

Dean and Tara are photographers. They take pictures of seniors for high school yearbooks. They charge a standard fee for taking the pictures.There also a charge for each set of prints ordered.So the number of sets of prints determines the total cost. Let s stand for the set of prints ordered.Write the rule that describes a students total cost. Print(s):0,1,2,3 Total cost :$20,$21,$22,$23

OpenStudy (anonymous):

Your base cost (for only taking pictures) will be $20 because 0 prints =$20 dollars. So one print costs $21-$20=__

OpenStudy (calculusxy):

we can use \(y = mx + b\). to find out m we can use \(\large \frac{y_2 - y_1}{x_2 - x_2}\). and then find out what b should be (which in this case it should be 20 because 0 sets of print equals to $20)

OpenStudy (calculusxy):

you would just need to find out what m is and then put it along with b (which you already know)

OpenStudy (anonymous):

they sections called functions though??

OpenStudy (anonymous):

so how could i use y=mx+b

OpenStudy (calculusxy):

y = mx + b <--- linear function but you can try to figure out from a quadratic

OpenStudy (anonymous):

so 20x+s

OpenStudy (calculusxy):

incorrect. you are only supposed to have one variable on the right side of the linear equation.

OpenStudy (calculusxy):

for example, a sample linear equation would be \(y = 6x + 3\)

OpenStudy (calculusxy):

on the right, you have only one variable x and two other integers (6 being the coefficient, or m, and 3 being the constant, or b)

OpenStudy (calculusxy):

did you figure out the slope (m) from the formula I gave you. you can use the two coordinates (0,20) and (1,21). i will put the equation again: \(\huge m = \frac{y_2 - y_1}{x_2-x_1}\)

OpenStudy (anonymous):

yes i understand i just cant find the rule to my question

OpenStudy (calculusxy):

can you give this a try at least?

OpenStudy (anonymous):

well the thing i am working on can only be answered once.

OpenStudy (calculusxy):

i am trying to give you the steps, but you are not following it

OpenStudy (anonymous):

ok so i got 1 for ur equation \